OpenStudy (anonymous):

helpp mee pleaseee someonee i dont get this... 2 square root of 27x^5y

OpenStudy (anonymous):

\[\sqrt[2]{27x ^{5}y}\]

OpenStudy (amistre64):

sqrt(3^2*3, x^2,x^2,x,y) 3x^2 sqr(3xy)

OpenStudy (anonymous):

so where does the 2 go?

OpenStudy (anonymous):

\[\sqrt{27x^5y}=\sqrt[2]{27x^5y}\] \[\sqrt[a]{b^c}= b ^{\frac{c}{a}}\]

OpenStudy (anonymous):

do you understand it now?